# Continuously tunable uniaxial strain control of van der Waals heterostructure devices

### v1

@@ Appendix B Strain transmission calibration
Figure 6: (a) Configuration for measuring the strain transmission on nanoflakes with a single piezo stack technique. The Si/SiO2 substrate and commercial SG were glued to top and bottom of the piezo stack with epoxy, respectively. The nanoflakes were transferred onto the SiO2 surface with standard dry-transfer procedures and the gold SGs were deposited on the single- and twisted-bilayer nanoflakes, and SiO2 surface. The green arrow presents the poling direction of the piezo stack. (b) Gauge factor of the gold SG as a function of the resistivity at 2 K. The different red symbols refer to different measurements. The purple hexagram indicates the gold SG in Fig. 3. All SGs here were directly deposited on the silicon substrate. The data can be fitted by a quadratic curve (dashed line). Our data is consistent with literature. Li, Hesketh, and Maclay (1994) (c) Temperature dependence of strain transmission from silicon to single hBN flake at different thickness. The thickness was measured by an atomic force microscope (Bruker Dimension Edge). (d) Similar to (c) measured on a randomly twisted hBN flake. Inset: the twisted hBN flakes with one gold SG on the bottom single layer region (pink square) and the other one on the twisted hBN region (blue square).
@@
To calibrate the gold SGs and to investigate the strain transmission through the vdW heterostructures, we have conducted measurements on multiple gold SGs using single piezo stack technique. Shayegan et al. (2003) The gold SGs were fabricated either directly on top of a Si/SiO2 substrate or above several single- or twisted-hBN flakes, as shown in Fig. 6(a). The silicon wafer is glued to one side of the piezo stack, while a commercial SG is affixed on the other side. At low temperature, the strain can be almost fully transferred to the silicon wafer Chu et al. (2012) which can be precisely measured by the commercial SG. After determining the strain of the silicon wafer, we calibrate the gold SGs by measuring the change of resistance of the gold SGs fabricated on silicon wafer, and obtain the gauge factor using Eq. 1. In Fig. 6(b), we plot the obtained gauge factor as a function of the resistivity of the gold SGs. The gauge factor ranges from 2.1 to 3.3, in agreement with the previous report,Li, Hesketh, and Maclay (1994) and follows a quadratic behavior as a function of resistivity. The fitted quadratic function allows us to determine the gauge factor of gold SGs in other experiments.
To further investigate the strain transmission through the vdW heterostructures, we deposited gold SGs on hBN flakes to estimate the strain transmission on nanoflakes with different thickness. As discussed above, the gauge factor was determined using the fitted quadratic function, as shown in Fig. 6(b). By comparing the strain measured from gold SGs both on Si wafer and on hBN flakes, we calculate the transmission factor as a function of temperature for both single and twisted hBN flakes in Fig. 6(c, d). We found that strain transmissions for single hBN flakes are consistently above 90%, and remain nearly independent of temperature and thickness below 20 K. The slightly reduced strain transmission atop of the twisted hBN suggests a strain loss, likely due to the incommensurate alignment between the twisted flakes. However, it remains fairly high and remarkably consistent with the ratio obtained from the Raman spectra in Fig. 3. Hence, depositing gold SG on nanoflake proves to be an effective method for locally determining the strain and its transmission when Raman spectroscopy is not feasible.

### v2

@@ Appendix B Gold strain gauge calibration
Figure 6: (a) Configuration for measuring the gold strain gauges with a single piezo stack technique. The Si/SiO2 substrate and commercial SG were glued to top and bottom of the piezo stack with epoxy, respectively. The gold SGs were directly deposited on SiO2 surface. The green arrow presents the poling direction of the piezo stack. (b) Gauge factor of the gold SG as a function of the resistivity at 2 K. The different red symbols refer to different measurements. The purple hexagram indicates the gold SG in Fig. 3(e). The data can be fitted by a quadratic curve (dashed line): $\rm GF=-0.0113\rho^{2}+0.2753\rho+1.553$
@@
To calibrate the gold SGs, we have conducted measurements on multiple gold SGs using a single piezo stack technique. Shayegan et al. (2003) The gold SGs were fabricated either directly on top of a Si/SiO2 substrate or above several single-hBN flakes, as shown in Fig. 6(a). The silicon wafer is glued to one side of the piezo stack, while a commercial SG is affixed on the other side. At low temperature, the strain can be almost fully transferred to the silicon wafer which can be precisely measured by the commercial SG.Chu et al. (2012) After determining the strain of the silicon wafer, we calibrate the gold SGs by measuring the change of resistance of the gold SGs fabricated on silicon wafer and obtain the gauge factor using Eq. (1). In Fig. 6(b), we plot the obtained gauge factor as a function of the resistivity of the gold SGs. The gauge factor ranges from 2.1 to 3.3, in agreement with the previous report,Li, Hesketh, and Maclay (1994) and follows a quadratic behavior as a function of resistivity. The fitted quadratic function allows us to determine the gauge factor of gold SGs in other experiments. We have also estimated the strain transmission through the vdW heterostructures by taking the ratio between the strain measured from gold SGs on hBN flakes and those on the silicon wafer. The ratio is typically above 90% and remains nearly independent of temperature and thickness below 20 K. While the variation of the gauge factor between different gold SG does not allow us to directly interpret this ratio as the actual strain transmission factor, this is in broad agreement with the results of Raman measurement presented in Fig. 3 in the main text.